OpenStudy (firejay5):

For each problem, you must: 1. Define the variables 2. Write the function to be maximized or minimized. 3. Write a system of inequalities. 4. Graph and write the vertices of the feasible region. 5. Solve the problem and answer the question. Mr. Wood's construction class at CHSW is going to be building tool sheds this year and then selling them to the public to raise money for their program. They need 10 sheets of dry wall and 15 studs to build a small shed, and 15 sheets of dry wall and 45 studs to build a large shed. The principal told Mr. Wood's that he has 60 sheets of dry wall and 135 studs available for his students to use. If the school makes a $390 profit on a shed and $520 on a large shed, how many of each type of shed should the students build in order to maximize their profits?

hero (hero):

Okay, you should be able to solve this one on your own.

hero (hero):

I showed you step by step how to do the previous two. Now it's your turn.

OpenStudy (firejay5):

aww really come on

hero (hero):

Sorry to tell you the truth but I'm here to help students. I'll help you with the first couple problems. Then after that, it's your turn.

hero (hero):

That's the way I've always done it.

OpenStudy (firejay5):

I need help on the maximization

hero (hero):

0. Given Small Shed = 10 sheets of dry wall, 15 studs per Large Shed = 15 sheets of dry wall, 45 studs per Materials Available: 60 sheets of dry wall, 135 studs 1. Define variables x = no. of small shed built y = no. of large sheds built 2. Write Relevant Equations/Inequalities Drywall = 10x + 15y <= 60 Studs = 15x + 45y <= 135 Profit = 390x + 520y Solve systems of inequalities for x and y 10x + 15y <= 60 15x + 45y <= 135 5(2x + 3y) <= 5(12) 2x + 3y <= 12 15(x + 3y) = 15(9) x + 3y <= 9 3y <= 12 - 2x 3y <= 9 - x 0 <= 3 - x x <= 3 y <= 2 Therefore, you need to make at least 3 Small Sheds and 2 Large Sheds in order to gain max profit. Now, find out how much Profit the school made: P(x,y) = 390x + 520y P(x,y) = 390(3) + 520(2) P(x,y) = 1170 + 1040 P(x,y) = 2210 The school made $2210 (max)

hero (hero):

This was not an easy one. This question would stump many

OpenStudy (firejay5):

where's #"s 3 - 5

hero (hero):

Sorry. Once again, I provided the solution. You have to be able to decipher it

OpenStudy (firejay5):

okay